OpenStudy (anonymous):

In the following linear equation, if x is decreased by 1, how is y affected? y = 2x - 3 A. y will decrease by 2 B. y will decrease by 1 C. y will increase by 1 D. y will increase by 2

OpenStudy (anonymous):

try plugging in x = 2 and x=1 and calculate the 2 values of y

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

x=2 y = 2(2) - 3 = 1 x = 1 y = 2(2) - 3 = -1 decrease in y = 1 - (-) = 2 A

OpenStudy (anonymous):

* 1 - (-1) = 2

OpenStudy (anonymous):

recall that a double negative in a subtraction changes to an add